OpenStudy (anonymous):

(2x-1)^(2/3)=x^(1/3) Solve for x

OpenStudy (anonymous):

(2x^+1)(2/3)=1/3x^ (1 1/3X^+2/3=1/3x^)3 4x^+2=x^ 3x^=-2 x^=-2/3 Impossible? im pretty sure i screwed up

OpenStudy (anonymous):

The answer is square root of 2, square root or 3.

OpenStudy (anonymous):

(2x-1)^(2/3)=x^(1/3) Rearrange to \[\sqrt[3]{(2x-1)^{2}}=\sqrt[3]{x}\] where the a is a 3 Take both sides to third power to cancel third roots. Then expand \[(2x-1)^{2}\] into \[4x^{2}-4x+1\]=x Subtract x from both sides to have \[4x^{2}-5x+1\]=0 Quadratic equation it. You get 1 or .25. Verify by plugging it into original. It does work but I'm too lazy to write it all out.

OpenStudy (anonymous):

You're right I was looking at the solution to a different question. Thank you.